OpenStudy (anonymous):

exponential equations 2 ^ x-1 = 3

OpenStudy (australopithecus):

if you are trying to solve for x use logs

OpenStudy (anonymous):

\[2 ^{x-1} = 3\]

OpenStudy (anonymous):

how

OpenStudy (australopithecus):

Use this rule: \[\log_a(a^x) = x\]

OpenStudy (australopithecus):

Log(x) = y is the same as x = 10^(y)

OpenStudy (anonymous):

could you demonstrate please

OpenStudy (australopithecus):

\[\log_{10}(x) = \log(x)\] just note that when you see log with no base it means log base 10

OpenStudy (anonymous):

how about in my problem ?

OpenStudy (anonymous):

if you can use my problem as an example i could figure out the rest of my questions

OpenStudy (jdoe0001):

\(\bf 2 ^{x-1} = 3\\ \quad \\ \textit{log cancellation rule of }{\color{blue}{ log_aa^x=x}}\\ \quad \\ log_2(2 ^{x-1}) = log_23\implies x-1=log_23\)

OpenStudy (australopithecus):

sure, 55^x = 55 \[\log_{55}(55^x) = \log_{55}(55)\] \[x = \log_{55}(55)\]

OpenStudy (australopithecus):

x = 1

OpenStudy (jdoe0001):

the log cancellation rule idea being, the log base, and the base of the inner quantity, being equal

OpenStudy (anonymous):

thank you

OpenStudy (jdoe0001):

\(\bf \textit{log cancellation rule of }{\large \color{blue}{ log_{\color{red}{ a}}{\color{red}{ a}}^x=x}}\)

OpenStudy (agent0smith):

Take log base 10, not log base 2... you can't solve log base 2 of 3 without using change of base formula

OpenStudy (australopithecus):

Just to note, x = 1 because, \[\log_{55}(55^1) = \log_{55}(55) = 1\]

OpenStudy (agent0smith):

Or take natural logs of both sides.

OpenStudy (agent0smith):

natural logs of both sides\[\Large \ln 2 ^{x-1} = \ln 3\]bring down the exponent\[\Large (x-1)* \ln 2 = \ln 3\]then divide both sides by ln 2 and then just add 1 to both sides

OpenStudy (australopithecus):

oh yeah that rule, \[\ln(a^x) = (x)\ln(a)\] Note, \[\ln_e(x) = \ln(x)\] e is Euler's number get use to seeing it

OpenStudy (australopithecus):

oops, made a mistake, \[\log_e(x) = \ln(x)\]

OpenStudy (australopithecus):

PLEASE IGNORE the \[\ln_e(x) = \ln(x)\] that is incorrect

OpenStudy (australopithecus):

well actually maybe not, to be honest not to sure if that notation is accepted, probably not

OpenStudy (australopithecus):

though